### v3

# SpikeSMOKE: Spiking Neural Networks for Monocular 3D Object Detection with Cross-Scale Gated Coding

Abstract
With the wide application of 3D object detection in some fields such as autonomous driving, its energy consumption is constantly increasing, making the research on low-power consumption alternatives a key research area. The spiking neural networks (SNNs), possessing low-power consumption characteristics, offer a novel solution for this research. Consequently, we apply SNNs to monocular 3D object detection and propose the SpikeSMOKE architecture, which represents a new attempt at low-power monocular 3D object detection. It’s well known that the discrete signals of SNNs can lead to information loss compared to artificial neural networks (ANNs), which limits their feature representation capabilities. To solve this problem, inspired by the synaptic filtering mechanism of biological neurons, we propose a new Cross-Scale Gating Coding Mechanism (CSGC), which can enhance feature representation by combining cross-scale fusion of attentional methods and gated filtering mechanisms. In addition, to reduce the computation and accelerate training, we present a novel light-weight residual block that can maintain spiking computing paradigm and the highest possible detection performance. Our method is effective on the KITTI, NuScenes-mini and CIFAR10/100 datasets. Compared to the baseline SpikeSMOKE under the 3D Object Detection, the proposed SpikeSMOKE with CSGC can achieve 11.78 (+2.82, Easy), 10.69 (+3.2, Moderate), and 10.48 (+3.17, Hard) on the KITTI autonomous driving dataset by $AP|_{R_{11}}$ at 0.7 IoU threshold, respectively. It is worth noting that the results of SpikeSMOKE can significantly reduce energy consumption compared with the results of SMOKE. And SpikeSMOKE-L (lightweight) can further reduce the amount of parameters by 3 times and computation by 10 times compared to SMOKE.
Index Terms:
Spiking Neural Networks, Monocular 3D Object Detection, Gating Coding, Lightweighting.
I Introduction
Although ANNs have shown excellent performance in various fields including computer vision (CV)[22], speech recognition (SR)[11], natural language processing (NLP)[19], etc, and have made great contributions to the development of artificial intelligence, they are facing with skyrocketing problems of energy consumption, with the increasingly complex models and progressively large volumes of data[56]. Especially in 3D object detection, the maximum computational complexity observed in existing algorithms has already reached 50 GMAC[28]. This not only increases hardware costs but also limits their applications in mobile devices and edge computing scenarios. To solve this problem, some researchers are turning to brain-inspired SNNs, because they have excellent characteristics of event-driven computation[61], biologically interpretable[15], and asynchronous temporal processing. These features enable SNNs to significantly reduce energy consumption and improve computational efficiency when dealing with sparse data and temporal dynamic signals.
At present, there has been a great deal of excellent research on SNNs applied to computer vision (CV), natural language processing (NLP) and other fields[53][50][35][8][58][16]. In terms of the evolution of network architecture, the early Spiking ResNet[57] laid the foundation for event-driven network architectures. Subsequently, EMS-ResNet[9] optimized this foundation, and MS-ResNet[16] further enhanced the network’s ability to process complex data, bringing new ideas to the field of computer vision. In the realm of natural language processing, Spikformer[60], which is the initial utilization of Transformer in SNNs, opened a new path for sequence modeling. The subsequent Spike-Driven Transformer v1/v2[49][48] continuously optimized, which not only offers robust sequence modeling capabilities, but also improves the efficiency of parallel processing of intricate language tasks. For object detection, current research focuses on 2D object detection using 2D image data or 3D object detection using 3D data. For example, in Spiking-PointNet [38], point clouds—which represent 3D spatial information—were utilized for the purpose of detecting 3D objects, marking the first successful application of point clouds within SNNs. In aspect of 2D image data, SpikeYOLO[31] has put forward a spike-driven object detection framework, which is capable of effectively extracting target features from 2D images and carrying out target localization and classification. However, SNNs still face some challenges in the field of object detection, especially in monocular 3D object detection. Since monocular 3D object detection requires processing additional depth information, traditional 2D object detection methods cannot be directly applied to 3D object detection. Up to now, there has been no mature exploration of monocular 3D object detection via SNNs, and the low-power characteristic of SNNs has not been utilized in this task.
Monocular 3D object detection combines 2D image information and depth estimation to achieve detection and localization of 3D objects, which has been widely used in the field of autonomous driving[27][47][10][18]. Currently, there are many excellent models in the field of ANNs. For example, MonoWAD [33] introduced an innovative weather-adaptive diffusion mechanism, which can effectively deal with changes in image features under different weather conditions, thereby enhancing the robustness and accuracy of monocular 3D object detection. LabelDistill[21] presented a label-guided cross-modal knowledge distillation method, which uses data from other modalities to assist in 3D object detection from 2D images. Mono3DVG[52] offered a fresh perspective by detecting 3D objects through detailed descriptions of visual appearance, and is able to accurately identify and locate objects with complex appearances. MonoCD[46], a network architecture based on auxiliary information, further optimized the performance of monocular 3D object detection. However, when applying monocular 3D object detection to resource-constrained edge devices like autonomous driving[25], its practical deployment will be restricted. These edge devices typically have limited computing resources and energy supplies, while existing ANN models often require high computational power and energy consumption to implement complex detection algorithms. Therefore, it is very important to implement a network model with low-power characteristics for monocular 3D object detection. Against this backdrop, brain-inspired SNNs models have attracted attention for their unique low-power characteristics, which give them potential advantages in handling monocular 3D object detection tasks and offer new ideas for solving this problem.
Based on the above analysis, we apply brain-inspired Spiking Neural Networks (SNNs) with low-power characteristics to monocular 3D object detection to construct a novel architecture SpikeSMOKE. In order to shrink the gap between the discrete signals of SNNs and the continuous signals of ANNs for feature expressiveness, we propose a cross-scale gated coding mechanism (CSGC) inspired by the filtering mechanism of biological neuronal synapses, which can reduce information loss and improve the detection performance of the model by fusing multi-scale attention. To further reduce the model’s power consumption while maintaining the spike computation paradigm, we design innovative light-weight residual blocks that can significantly reduce the computational load and energy consumption while maintaining the network’s depth and complexity. To verify the effectiveness of the proposed method in this paper, we conducted a large number of experiments, which cover the widely used KITTI and NuScenes-mini datasets in the field of autonomous driving, as well as the CIFAR10 and CIFAR100 datasets in the field of computer vision. The contributions of this paper can be summarized as follows.
(1) To enhance the energy efficiency of monocular 3D object detection, we propose a spiking architecture SpikeSMOKE based on SMOKE leveraging the low power consumption feature of brain-like SNNs.
(2) To improve the information representation capability of SpikeSMOKE, we design a novel parallel cross-scale gated coding mechanism CSGC based on attention. Additionally, we propose a light-weight residual block integrated into the SpikeSMOKE model.
(3) Compared to the baseline SpikeSMOKE, our enhanced model with CSGC achieves superior performance on the KITTI dataset. We further validate the generalizability of the proposed CSGC encoding strategy through experiments on the NuScenes-mini and CIFAR-10/100 datasets. Meanwhile, compared with SMOKE, SpikeSMOKE significantly reduces energy consumption in 3D object detection.
This paper is organized as follows. In section II, we describe the kinetic formulation of LIF for spike neurons, the model architecture for monocular 3D object detection, the coding mechanism, and lightweighting methods. In section III, we describe the macroscopic architecture of SpikeSMOKE, the detailed implementation of cross-scale gated coding (CSGC) and lightweighted residual structures. In the IV section, we present the experimental results of our methods and the results of ablation experiments. In the V section, we summarize the paper.
II Related works
II-A Spike Neuron
Spike neurons simulate the behavior of biological neurons transmitting signal sequences through synapses[43], transmitting temporal signals and spatial information through the membrane voltage U. This mechanism enables spike neurons to efficiently process dynamic information while maintaining low energy consumption, which is particularly significant when dealing with sparse data and time-series tasks. The commonly used neurons in spiking neural networks consist of integrate-and-fire (IF)[32] neurons and leaky integrate-and-fire (LIF) [40][17]neurons. These neuron models simulate the signal processing of biological neurons through different mechanisms[24]. We used LIF neurons and described their neuronal dynamics formulation as follows:
$U_{i}^{l}[t]=H_{i}^{l}[t-1]+\sum_{j}W_{ij}S_{j}^{l}[t]$ (1)
[…]
$H_{i}^{l}[t]=\tau U_{i}^{l}[t](1-S_{i}^{l}[t])+U_{reset}S_{i}^{l}[t]$
[…]
II-B Monocular 3D Object Detection
Monocular 3D object detection relies solely on the 2D image information obtained from a monocular camera to detect and locate 3D objects in a scene. However, the lack of depth information in monocular images makes it highly challenging to recover 3D information from 2D images. To address this challenge, researchers have proposed a variety of methods for monocular 3D object detection, which can generally be divided into three categories: two-stage methods, single-stage anchor-based methods, and single-stage anchor-free methods. Two-stage methods, such as MonoLSS[26], typically first generate a series of candidate regions and then perform fine classification and regression on these regions to determine the final 3D object bounding boxes. This method is prone to introducing 2D noise[30], which to some extent affects the accuracy and reliability of detection. Single-stage anchor-based methods, such as GrooMeD-NMS[23], define a series of anchor points on the image in advance and then perform classification and regression on these anchor points to directly generate 3D bounding boxes. These methods usually require complex data preprocessing and non-maximum suppression[6][36]. Compared with the first two categories of methods, single-stage anchor-free methods (SMOKE[30], MonoCD [46], MonoDGP [34]) have the significant advantages of a simple model and high computational efficiency, making them more suitable for real-time operation on resource-constrained edge devices.
In this paper, we use the SMOKE [30] architecture of Single-stage anchor-free methods as a base architecture, which both ignores the redundancy associated with 2D detection frameworks and does not require additional data. The SMOKE architecture continues the keypoint detection of centernet [7], which discards the traditional 2D detection module and retains only the 3D detection part, and improves the parameter convergence and detection accuracy by multi-step disentanglement. Specifically, it decomposes the attribute regression of the 3D bounding box into multiple steps, optimizing each parameter progressively. The 3D bounding box is obtained by predicting the object’s 3D projection center on the image plane along with attribute variables. The attribute regression of the 3D bounding box is decoupled into an 8-tuple $(\alpha_{x},\alpha_{y},\alpha_{z},\alpha_{l},\alpha_{w},\alpha_{h},sin\beta,cos\beta)$. These parameters are optimized by a loss function, and then transformed into the actual parameters for constructing the 3D bounding box, including the object’s center position $(x,y,z)$, dimensions $(w,h,l)$, and orientation angle $(\theta)$.
II-C Coding Mechanism
In a spiking neural network, data encoding is a crucial step for efficient information processing. Initially, input data must be transformed into spike sequences for further processing by neurons, a process that involves selecting from several encoding methods. Rating coding [42] encodes information based on the average rate of neuron firing, the higher the firing rate of neurons, the greater the signal strength or information content they represent. Its advantage lies in its simple implementation and similarity to the behavior of biological neurons. However, it requires a longer time window to accurately estimate the firing rate, which may lead to information delay. Phase coding [20] expresses information by the temporal position of spikes relative to a reference event. For instance, in a periodic stimulus, the phase of a spike can precisely indicate specific signal features. Temporal coding [59] represents information based on the time intervals and order of spikes, effectively utilizing the temporal characteristics of spikes to handle dynamic and sequential data. Frequency coding involves the frequency at which neurons emit spikes over a specific time period. It represents information through the repetition rate of spikes and emphasizes short-term frequency changes rather than long-term average firing rates. In addition to this, there are various complex coding methods ([37][45] ), and choosing an appropriate coding method is crucial for the representation performance of the model. In order to reduce the gap between the discrete and continuous signals in the spiking neural network, we propose a novel gated coding method in conjunction with the cross-scale fusion unit. This encoding method introduces a gating mechanism that more effectively controls the flow and processing of information.
Figure 1: SpikeSMOKE overall network architecture. The input is an image and the output is the 3D bounding box of the object detection and the object category. A CSGC coding mechanism is introduced, based on a cross-scale attention fusion, to reduce the performance loss due to data transformation and enhance the representation of complex features.
II-D Light-Weight Model
Light-weight model refers to reducing the number of parameters and computation and compressing the model training time by removing or transforming part of the model architecture and connections[44][51]. There are several common methods for lightweighting models in neural networks, a) Pruning: Chen et al [4] proposed an SNN pruning regeneration mechanism based on neuron criticality to remove unnecessary connections or neurons in the network. This mechanism can identify and remove connections or neurons that contribute less to the model’s performance, thereby simplifying the network structure and reducing unnecessary computations. SSNN [2] designed sparse RSNN by pruning randomly initialized model. b) Model quantization: Model quantization reduces the model’s storage requirements and computational load significantly by converting the model’s weights and activations from floating-point numbers to low-bit representations, such as binary or ternary quantization. Spiking-Diffusion [29] proposed a diffusion model based on vector quantization discretization. c) Model distillation: TSSD [62] proposed a spatio-temporal distillation method, which transfers the knowledge from a complex large model (the teacher model) to a simpler small model (the student model). This enables the student model to inherit the performance of the teacher model while maintaining low computational complexity. d) Depth-wise separable convolutions: Depth-wise separable convolutions decompose standard convolutions into two steps: depth-wise convolutions and point-wise convolutions, significantly reducing the computational load of convolution operations. The first proposed in MobileNets [14], and later extended and applied widely in the field of ANN ( [5], [3]). There has been a great deal of experimental or theoretical evidence for the effectiveness of depth-wise separable convolutions, and our proposed lightweight residual unit originates from this idea, aiming to further optimize the model’s computational efficiency and performance.
III Methods
III-A SpikeSMOKE
The SpikeSMOKE architecture is shown in Figure 1. The design of this architecture is based on the single-stage object detection network SMOKE, which has the advantages of ignoring the redundancy of 2D detection frameworks and not requiring additional data. We choose this architecture because it can simplify the complexity of the model while maintaining efficient detection performance.
Backbone: The architecture DLA34 can take advantage of fusing feature maps of different scales through the depth convergence feature. To fully transform the model into a SNNs-based architecture, we introduced the spiking rate of spiking neurons to replace the ReLU activation function in DLA34, which is very crucial. Meanwhile, to enhance the processing capability of spike signals, we have proposed the CSGC module in backbone, including channel attention mechanism and spatial attention mechanism. A detailed introduction to CSGC is in Section Methods B.
Neck: The design of the neck aims to further optimize the transmission and processing of features, ensuring that the model is fully driven by spikes. In the DLAup structure, we ensure that a spiking neuron is used before each convolution operation to convert the continuous signals into discrete spike signals, thereby maintaining the spike-driven characteristics. Meanwhile, deformable convolutions were employed to further enhance feature extraction.
Head: This component consists two branches. One is keypoint classification used by heatmap, which is mainly responsible for analyzing the input feature maps to identify the locations of object keypoints and generate heatmaps to represent the distribution of these keypoints. The other is 3D bounding box regression, which is in charge of predicting the 3D bounding box parameters of each detected object based on the information in the feature maps, including the center coordinates, size, and rotation angles of the bounding boxes. This two branches can process the feature maps from the Neck network to obtain the 3D object detection results.
Figure 2: Gated coding is different from direct coding. The proposed gated coding mechanism utilizes a cross-scale fusion module that incorporates a gating unit (CSGC) to mimic the filtering mechanism of biological neuronal synapses to produce a filtering effect on the input feature map.
III-B Cross-Scale Gated Coding (CSGC)
When converting ANN to SNN, the fundamental contradiction of sparsity and discreteness of SNN is faced, which weakens the modeling ability of continuous spatial structures. Especially in monocular 3D detection tasks, the model needs to recover nonlinear 3D geometric information from 2D images, which poses high requirements for the multi-scale context awareness ability and key structure retention ability of features. However, standard SNN encoders (such as direct ANN2SNN conversion) lose a large amount of information during the spiking process. Therefore, in order to enhance the feature expression ability and reduce information loss, we propose a cross-scale fusion of attention coding unit named as CSGC, as shown in Figure 2.
For CSGC, we design a parallel architecture including channel attention and spatial attention and a gate control unit. The Channel Attention addresses the issue of which feature channels are more important for 3D geometric reasoning. The Spatial Attention is used to focus on which areas in an image are more critical to the current target scale. The original input tensor $x\in{R^{C\times H\times W}}$ is firstly repeated at each time step to obtain information in the time dimension. Then, the original tensor is fed into both the channel attention and spatial attention modules. In the channel attention part, we use a Linear-ReLU-Linear structure to learn and update the weights in the channel dimension by the Equation (2):
$CA(x)=Linear\left(ReLU(Linear(x))\right).$ (2)
The number of neurons in the middle layer of the network is set to the number of input channels /ratio, and the ratio is set to 4. This structure can effectively highlight important channel features, thereby enhancing the model’s ability to perceive key information. In the spatial attention part, we conduct feature extraction on the same feature map by different convolution kernel (Equation (3)). Small convolution kernels are employed to capture local subtle features for small objects, enabling the detection of small-sized objects. Conversely, larger convolution kernels are utilized to cover larger areas for detecting large objects and effectively capturing global information. To better integrate information from different scales, we employ three different sizes of convolution kernels and assign dynamic weights to the feature maps after different convolutional processing by the learnable parameters $\alpha$, $\beta$ and $\gamma$, respectively, with the initial weights all set to 0.5. At the same time, we invoke the residual idea of [12] to connect the original feature map with the output part. This not only enhances the features but also prevents information loss, ensuring that the model can fully utilize all useful information from the original input.
$SA(x)=$ $Conv_{3\times 3}\left(ReLU\left(Conv_{3\times 3}(x)\right)\right)\cdot\alpha+$ (3)
$Conv_{5\times 5}\left(ReLU\left(Conv_{5\times 5}(x)\right)\right)\cdot\beta+$
$Conv_{7\times 7}\left(ReLU\left(Conv_{7\times 7}(x)\right)\right)\cdot\gamma+$
$x.$
The output spike information after the above channel and spatial attention processing is controlled and adjusted by a Sigmoid gating unit, which is defined as Equation (4):
$O(x)=\sigma\left(SA(x)\odot CA(x)\right),$ (4)
where $\sigma$ denotes the Sigmoid function, which maps values to the range (0,1), and $\odot$ denotes the Hadamard product. Then, it performs the Hadamard product operation with the output of the spiking neuron, imitating the filtering function of the synapses, as defined by Equation (5):
$\widehat{O}(x)=O(x)\odot SNN(x),$ (5)
where $O(x)$ represents the attention scores obtained after the parallel channel attention and spatial attention modules, indicating the importance of the detected features. $SNN(x)$ denotes the output spikes from the original feature x after processing through the Leaky Integrate-and-Fire (LIF) neuron. Since in SNNS, features exist in the form of binary pulses, direct weighting would undermine their sparsity and event-driven characteristics. Therefore, CSGC takes the attention score $O(x)$ as the gate signal and performs a Hadamard product with the binary output (0 or 1) of $SNN(x)$, which is generated by an LIF neuron, allowing only the spikes in the high-attention region to pass through, thereby obtaining the final output $\widehat{O}(x)$. This essentially simulates the filtering mechanism in biological synapses, achieving feature screening while maintaining the sparsity of spikes.
III-C Light-weight Residual
Figure 3: Light-weight residual blocks. The rightmost figure uses depth-wise convolution and point-wise convolution to replace vanilla convolution.
When SNN is applied to high-complexity tasks such as monocular 3D object detection, the residuals of standard convolution need to be repeatedly executed over multiple time steps, and the fully connected channel interaction can lead to a large amount of invalid computations, which goes against the original intention of energy efficiency of the SNN. Therefore, we introduce a novel light-weight residual block, as shown in Figure 3, is introduced to significantly reduce the number of model parameters and computational complexity by integrating depth-wise separable convolution technology and a membrane-based shortcut method, thereby enhancing the efficiency of model training. In terms of implementation details, the lightweight residual block first employs depth-wise convolution to perform independent convolution operations on each channel of the input data. Compared to traditional convolution, depth-wise convolution has the advantage of operating on each channel individually, thus avoiding redundant computations across channels and significantly improving computational efficiency. After completing the depth-wise convolution, we then apply point-wise convolution , with the purpose of integrating information from various channels and enhancing the expressive power of features. Then, we apply depth-wise convolution again to further extract and strengthen features. During this process, to maintain the flow of information and prevent the vanishing gradient problem, we introduce a shortcut path based on the membrane method
For regular membrane-based residual block, the computational volume $\alpha_{1}$ and number of parameters $\beta_{1}$ can be calculated by expression (6) and the expression (7) respectively:
$\alpha_{1}=k\cdot k\cdot C_{in}\cdot C_{out}\cdot W_{out}\cdot H_{out},$ (6)
$\beta_{1}=k\cdot k\cdot C_{in}\cdot C_{out}.$ (7)
When the above light-weight residual block is introduced, the computational volume $\alpha_{2}$ and parametric quantities $\beta_{2}$ can be calculated by the expression (8) and the expression (9) respectively:
$\alpha_{2}=k\cdot k\cdot C_{in}\cdot W_{out}\cdot H_{out}+C_{in}\cdot C_{out}\cdot W_{out}\cdot H_{out},$ (8)
$\beta_{2}=k\cdot k\cdot C_{in}+C_{in}\cdot C_{out}.$ (9)
For the SpikeSMOKE with light-weight residual block, the convolution kernel size $k$ is $3\times 3$, $C_{\text{out}}=2C_{\text{in}}$, $W_{\text{out}}=\frac{1}{2}W_{\text{in}}$, $H_{\text{out}}=\frac{1}{2}H_{\text{in}}$. Then, the ratio of the computational effort can be calculated by the Equation (10):
$\frac{C_{in}\cdot W_{out}\cdot H_{out}(3k^{2}+2C_{in})}{3k^{2}\cdot C_{in}\cdot C_{out}\cdot W_{out}\cdot H_{out}}=\frac{1}{2C_{in}}+\frac{1}{27}.$ (10)
Similarly, the ratio of their parametric quantities is the Equation (11):
$\frac{k^{2}\cdot C_{in}+C_{in}\cdot C_{out}+k^{2}C_{out}}{k^{2}C_{in}\cdot C_{out}+k^{2}C_{out}\cdot C_{out}}=\frac{1}{2C_{in}}+\frac{1}{27}.$ (11)
Observably, compared to the pre-improvement phase, the proposed light-weight residual block diminishes computational and parametric quantities by approximately 27 times when $C_{in}$ is large enough. The exact reduction ratio depends on the $C_{in}$. This novel lightweight residual block, through carefully designed convolution operations and shortcut connections, achieves a reduction in computational load and parameter count while preserving model performance as much as possible, offering new perspectives for the optimization of deep learning models.
IV Experiments
IV-A Datasets
KITTI.
The KITTI dataset is widely used in the field of autonomous driving and computer vision. The monocular 3D object detection utilizes single-camera data for stereo object detection across the Car, Pedestrian, and Cyclist categories, with 3,712 pictures in the training set and 3,769 pictures in the validation set, encompassing diverse scenarios such as city roads and highways. The detected objects include 3D bounding boxes, bird’s eye view bounding boxes, 2D bounding boxes, and can be classified into Easy, Moderate, and Hard categories based on detection complexity. We used an evaluation metric called 11-point Interpolated Average Precision metric [39], which approximates the shape of the Precision/Recall curve, at each difficulty level. It is defined as the Equation (12):
$AP|N=\frac{1}{|N|}\sum_{n\in N}f_{inter}(n)$ (12)
Where N denotes the set of recall levels that represent the exact intervals, this experiment takes $R_{11}=\{\frac{1}{10},\frac{2}{10},\frac{3}{10},\cdots,1\}$. $f_{inter}(n)$ denotes the precision obtained after interpolation calculation, instead of averaging the precision values, the maximum precision greater than or equal to the current recall is taken in calculating the precision. It is defined as the Equation (13):
$f_{inter}(n)=\max_{n^{\prime}:n^{\prime}\geq n}f(n^{\prime})$ (13)
TABLE I: The results of 3D object detection and detection under BEV are performed on
the validation set of KITTI dataset for the car class.
Methods(IoU) Parameters(M) Power(pJ) 3D Object Detection Birds’ Eye View Runtime(s)
Easy Moderate Hard Easy Moderate Hard
ANN SMOKE(0.7) 19.51 2.17E+11 14.76 12.85 11.5 19.99 15.61 15.28 0.03
SMOKE*(0.5) 19.51 2.17E+11 29.16 25.22 21.47 32.52 29.59 25.86 0.03
SMOKE*(0.7) 19.51 2.17E+11 12.03 11.14 10.92 17.97 13.08 12.06 0.03
FCOS3D(0.7) 50.62 3.04E+11 13.9 11.61 10.98 18.16 14.02 13.85
MonoDLE(0.7) 20.82 2.87E+11 17.45 13.66 11.68 24.97 19.33 17.01 0.04
MonoDTR(0.7) 54.25 5.42E+11 24.52 18.57 15.51 33.33 25.35 21.68 0.037
MonoLSS(0.7) 21.51 2.81E+11 25.91 18.29 15.94 34.70 25.36 21.84 0.035
SNN SpikeSMOKE(0.7) 19.51 5.97E+10 8.96 7.49 7.31 12.07 9.95 9.32 0.09
SpikeSMOKE-CSGC(0.7) 19.56 6.04E+10 11.78 10.69 10.48 15.67 13.68 11.83 0.105
SpikeSMOKE(0.5) 19.51 5.97E+10 20.87 19.72 16.89 27.64 22.47 21.65 0.09
SpikeSMOKE-CSGC(0.5) 19.56 6.04E+10 28.83 22.75 19.44 36.23 26.88 25.75 0.105
TABLE II: The results of 3D object detection and detection under BEV are performed on
the test set of KITTI dataset for the car class.
Methods
Parameters
Power
3D Object Detection Birds’ Eye View
Easy Moderate Hard Easy Moderate Hard
SMOKE-ANN* 19.51 2.17E+11 28.97 23.57 20.21 30.48 27.76 25.81
SpikeSMOKE 19.51 5.97E+10 21.80 15.23 14.99 25.99 20.69 17.92
SpikeSMOKE-L 6.32 2.24E+10 19.46 10.52 10.23 24.01 15.74 15.37
SpikeSMOKE -LCSGC
6.37 2.29E+10 20.64 15.32 12.93 25.15 18.98 18.08
The table is evaluated using the $AP|_{R_{11}}$ with a 0.5 IoU threshold.
TABLE III: The results of 2D object detection is performed on
the validation set of KITTI dataset for the car class.
Methods 2D Object Detection
Easy Moderate Hard
SMOKE-ANN* 80.49 72.01 68.76
SpikeSMOKE 73.32 62.85 55.67
SpikeSMOKE-CSGC 75.58 65.49 64.37
SpikeSMOKE-L 51.59 44.01 42.71
SpikeSMOKE-LCSGC 52.33 49.04 43.04
TABLE IV: Classification results on the classification task
dataset CIFAR-10.
Architecture
Coding
Schemes
Time
Steps
CIFAR10
Acc. (%)
ResNet-19[37] Phase Coding 8 91.40
VGG-16[37] Temporal Coding 100 92.68
ResNet-19[37] Rate Coding 6 93.16
MS-ResNet-18 Direct Coding 6 94.92
MS-ResNet-18 CSGC Coding 6 95.98(+1.06)
TABLE V: Classification results on the classification task
dataset CIFAR-100.
Methods Architecture
Spike
Params
Time
Steps
CIFAR100
Acc. (%)
ANN[37] MS-ResNet-18 $\times$ 12.54 N/A 80.67
MS-ResNet-SNN[37] MS-ResNet-18 $\checkmark$ 12.54 6 76.41
CSGC-SNN MS-ResNet-18 $\checkmark$ 12.72 6 79.58
MS-ResNet-18 $\checkmark$ 12.72 4 77.97
CIFAR-10/100.
The CIFAR-10 and CIFAR-100 datasets are commonly used for image classification. The CIFAR-10 dataset contains 10 categories of color images, each containing 6000 images of size 32x32 pixels, for a total of 60, 000 images. The CIFAR-100 dataset extends CIFAR-10 with 100 categories, each containing 600 images of size 32x32 pixels. contains 600 images, for a total of 60, 000 images with image size of 32x32 pixels.
NuScenes-mini.
The NuScenes-mini dataset is a subset of the NuScenes dataset, which is collected by 6 cameras, 1 Lidar, 5 millimetre wave radars, GPS and IMU, covering complex traffic conditions as well as variations in weather and light. The dataset contains 8 attributes of 3 types of objects, namely vehicles, bicycles, and pedestrians, and 10 scenes with continuous images of about 20 seconds each, which is an important dataset for 3D object detection in the field of autonomous driving. The dataset presents a composite metric, the nuScenes Detection Score (NDS), which was combined by calculating weighted sums (mAP, mATE, mASE, mAOE, mAVE, and mAAE).
IV-B Experiments Setup
We used random horizontal flip, random scale and shift as data augmentation aiming to increase the diversity of training data. In the network design, we set the number of groups for group normalization to 32 and for less than 32, we take 16. based on the analysis in literature [41], we set $\begin{bmatrix}\bar{h,}\bar{w,}\overline{l,}\end{bmatrix}^{\top}=[1.63,1.53,3.88]^{\top}and\begin{bmatrix}\mu_{z},\sigma_{z}\end{bmatrix}^{\top}=[28.01{\operatorname*{,16.32}}]^{\top}.$ The input image resolution size of the network is 1280x384, and after multiple downsampling, the size is reduced to 32 times of the original size. We trained 172 epoch using four NVIDIA GeForce RTX 4090 GPU, with a size of 4 per batch and a learning rate set to $1.25\times 10^{-4}$. The decay strategy for the learning rate was to reduce the learning rate to 10 times of the original at epoch 47 and 90. We used a threshold of 0.25 to filter the detection object. The lightweight strategy adopted in the experiment only acts on the backbone of feature extraction. Additionally, to more clearly present the configuration of each layer in the network, we have used a tabular format for detailed explanation, as shown in Table VIII. The table lists key information including the size of the convolutional kernels, the number of input and output channels, the type of neurons, and the dimensions of the input and output, all of which are essential parameters when constructing a deep learning model.
Figure 4: Ablation experiments in CSGC with different attention modules at different time steps.
TABLE VI: Results on the validation set of the NuScenes-mini dataset.
Methods mAP NDS
SMOKE 31.3 37.2
SpikeSMOKE 25.3 29.8
SpikeSMOKE-CSGC 26.6 31.2
Figure 5: The visualization results from monocular 3D object detection on the KITTI dataset present a clear and intuitive representation of the performance and accuracy of the detection algorithm in identifying and localizing 3D objects within the complex road scenarios depicted in the dataset.
TABLE VII: Ablation experiments on the threshold of spike neurons.
Methods vth=0.25 vth=0.5 vth=0.75 vth=1.0 2D Object Detection
Easy Moderate Hard
SpikeSMOKE-CSGC $\checkmark$ 64.31 53.84 51.62
$\checkmark$ 72.06 61.27 60.67
$\checkmark$ 75.58 65.49 64.37
$\checkmark$ 54.10 45.88 44.98
TABLE VIII: The detailed settings of the SpikeSMOKE-CSGC architecture.
Layer Name Input Size Layer Output Size
Base_layer (T,B,3,384,1280) 7×7Conv,32,stride1,padding3→GN→LIF (T,B,32,384,1280)
CSGC (T,B,32,384,1280) Channel-Attn + Spatial-Attn (3×3/5×5/7×7Conv) + Sigmoid (T,B,32,384,1280)
Avgpooling (T,B,32,384,1280) AvgPool2d(kernel=2, stride=2) (T,B,32,192,640)
Layer0 (T,B,32,192,640) 3×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,64,96,320)
Layer1 (T,B,64,96,320) 4×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,128,48,160)
Layer2 (T,B,128,48,160) 6×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,256,24,80)
Backbone Layer3 (T,B,256,24,80) 4×(LIF→3×3Conv→LIF→3×3Conv)-LIF (T,B,512,12,40)
Upblock0 LIF→3×3Conv(512-256)→LIF→ConvTranspose(2×)
Upblock1 LIF→3×3Conv(256-128)→LIF→ConvTranspose(2×)
Neck Upblock2 LIF→3×3Conv(128-64)→LIF→ConvTranspose(2×) (T,B,64,96,320)
class head (T,B,64,96,320) LIF→Conv3×3(64→256)→GN→LIF→Conv1×1(256→3) (B,3,96,320)
Head regression head (T,B,64,96,320) LIF→Conv3×3(64→256)→GN→LIF→Conv1×1(256→8) (B,8,96,320)
IV-C Main Results
Object Detection Performance on KITTI.
We validate the performance on the validation and test sets of the KITTI dataset for 3D object detection and bird’s-eye view under three levels of Easy, Moderate and Hard. * denotes the experimental results obtained by us after configuring the environment and training based on the SMOKE. The 3d object detection of our proposed SpikeSMOKE with CSGC (SpikeSMOKE-CSGC) can reach 28.83/11.78 (Easy), 22.75/10.69 (Moderate), 19.44/10.48 (Hard), and for the BEV detection, they can reach 36.23/15.67 (Easy), 26.88/13.68 (Moderate), 25.75/11.83 (Hard), as shown in Table I, where the metrics is evaluated by $AP|_{R_{11}}$ at 0.5/0.7 IoU threshold. From the results of these experiments, it can be found that although our detection performance has a little gap compared with the SMOKE-ANN*, energy consumption has been significantly reduced. For example, we calculate the energy consumption of the hard category for the 0.7 IoU threshold and find that it can be reduced by 72.2%, while the detection performance is reduced by only 4%. These experimental results are obtained because SMOKE-ANN* relies on continuous value activation and can precisely model 3D geometric attributes. However, SpikeSMOKE converts features into binary spike sequences (0/1), and information is inevitably lost during the encoding process. Meanwhile, the low-power consumption feature of our SNN event-driven mechanism reduces energy efficiency. The formula for calculating energy consumption is as follows:
[…]
where 0.9 denotes the consumption per accumulator operation and 4.6 denotes the multiplier word operation [54][13]. With a time step T=8, we achieved a runtime latency of 0.105s when performing inference on 4 NVIDIA GeForce RTX 4090 GPUs. Since we extended the time dimension when training on the GPU, it is reasonable that the latency would be larger compared to the traditional ANNs model.
It is well known that monocular 3D object detection is commonly used in resource-constrained scenarios such as edge devices, embedded devices, etc., therefore, lightweighting is an important part of research. Based on this, we further discuss SpikeSMOKE-CSGC’s lightweighting by light-weight residul block. The experiments show that the number of parameters is only 1/3 of SMOKE-ANN*, and the amount of computation is only 1/10, as shown in Table II. As a result, the SpikeSMOKE model may provide a new effective solution to reduce the energy consumption of monocular 3D object detection and improve its energy efficiency.
Comparing with the baseline SpikeSMOKE, the proposed SpikeSMOKE-CSGC shows significant improvements in both 3D object detection and BEV detection, with gains of 2.82, 3.2, and 3.17 for 3D object detection, and 3.6, 3.73, and 2.51 for BEV detection, as displayed in Table I. The results of the SpikeSMOKE-CSGC with lightweight treatment (SpikeSMOKE-LCSGC) can also demonstrate notable enhancements compared to that of the SpikeSMOKE-L, as indicated in Table II. Therefore, the improved CSGC method has a significant effect on 2D/3D object detection.
Efficiency and Generalizability Validation for CSGC.
We validate the classification results of the proposed CSGC coding strategy on MS-ResNet18 on the classification dataset CIFAR-10/100, as shown in Table IV and Table V. We note that the accuracy rate of the MS-ResNet18 with CSGC coding on the CIFAR-10 is 1.06% higher than the direct coding method. On the CIFAR-100, the proposed method MS-ResNet18-SNN with CSGC coding can achieve 79.58% by 6 time steps, which is 3.17% higher than MS-ResNet18-SNN. Meanwhile, we validated its effectiveness on the NuScenes-mini dataset, as shown in Table VI. Consequently, our proposed CSGC coding method is effective and generalized.
Ablation Studies on the Effects of Various Attentions.
We performed ablation experiments on 2D object detection about spatial attention (SA) and channel attention (CA) of CSGC at time steps of 4, 6, and 8, respectively, as shown in Figure 4. We observe that the detection performance is better than that of the baseline when SA or CA is used, which means that each module of CSGC is effective. Obviously, when CA and SA are used together, their detection results are better. Additionally, as the time step increases, the detection performance gets better.
Ablation Studies of Individual Neuronal Thresholds.
Since the threshold of the spike neurons has a great impact on the detection performance of the model, we conduct some ablation experiments on 2D detection by different neuron thresholds in order to get the most suitable LIF neuron threshold. Based on the experiment results, we obtain that the best performance is obtained when the neuron threshold vth=0.75, which is used in this paper, as shown in Table VII.
Qualitative Results.
We provide visualization results of the monocular 3D object detection on the KITTI dataset and compare them with the SpikeSMOKE and SMOKE, as shown in Figure 5. It can be observed that the detection performance of the Spikesmoke-CSGC architecture is superior to that of the SpikeSMOKE architecture and is also closer to that of the SMOKE architecture. Through these results, we can visually understand the model’s capability to accurately identify and locate 3D objects in complex road scenarios, thereby affirming the significant potential application of this technology in areas such as autonomous driving.
V Conclusion
With the widespread use of 3D object detection in applications such as autonomous driving, the low energy consumption problem is getting more and more attention. As widely acknowledged, low-power consumption stands out as a key feature of brain-like SNNs, offering a potential new solution for energy-efficient 3D object detection. Based on the SMOKE architecture and SNNs model, we have constructed a novel Spiking SMOKE architecture named SpikeSMOKE for monocular 3D object detection to reduce its energy consumption. In this architecture, the discrete signaling characteristics of SNNs may result in information loss and restrict their capacity for feature representation. Therefore, we have proposed a new cross-scale gating coding module named CSGC that can enhance the feature expression ability of the model, inspired by the synaptic filtering process in biological neurons. Furthermore, we have also proposed a lightweight residual block to reduce the computational effort while maintaining the impulse computation paradigm. We have conducted a large number of experiments on the KITTI, NuScenes-mini and CIFAR10/100 datasets to verify that the SpikeSMOKE architecture and CSGC module proposed in this paper are effective. The experimental results on the KITTI dataset, show that SpikeSMOKE can achieve higher energy efficiency compared to SMOKE, e.g., 72.2% reduction in energy consumption on Hard category, while the detection performance drops by only 4%. Moreover, the experimental results also show that the CSGC-based SpikeSMOKE can achieve significant improvement over the baseline SpikeSMOKE. SpikeSMOKE-L (lightweight) can further reduce the amount of parameters by 3 times and computation by 10 times compared to SMOKE. In the CIFAR-10/100 classification task, the CSGC encoding strategy improves the correctness by 1.06% and 3.17%, respectively, validating its generality. Overall, the SpikeSMOKE architecture and CSGC mechanism can provide an efficient and feasible solution for low-power monocular 3D object detection. In the future, we will continue to explore how to utilize the SNNs model to further enhance the performance of monocular 3D object detection and further reduce its energy consumption.